OpenStudy (anonymous):

show that \(\dfrac{n!}{(n-r)!r!}\) is integer , w/o saying "is an integer because this represents the "count" of number of ways of choosing "r" things from "n" things"

OpenStudy (thomas5267):

@ganeshie8 ........

ganeshie8 (ganeshie8):

lol this looks like an insult to my previous proof haha!

OpenStudy (anonymous):

na :|

OpenStudy (anonymous):

i really likes it xD

ganeshie8 (ganeshie8):

lets start with the inequality of greatest integer function since i know you're excited about legendre formula :) for any two real numbers \(a, b\) we have : \[\lfloor a+b \rfloor \ge \lfloor a\rfloor + \lfloor b \rfloor \]

OpenStudy (anonymous):

aha

ganeshie8 (ganeshie8):

for example \[\lfloor 2.6+0.6 \rfloor = 3 \ge \lfloor 2.6\rfloor + \lfloor 0.6 \rfloor = 2\]

OpenStudy (anonymous):

ok i agree

ganeshie8 (ganeshie8):

next consider \(r + (n-r)\) : \[\left\lfloor\dfrac{r+(n-r)}{p^k} \right\rfloor \ge \left\lfloor\dfrac{r}{p^k}\right\rfloor + \left\lfloor\dfrac{n-r}{p^k}\right\rfloor \] where \(p\) is a prime factor of \(r!(n-r)!\)

ganeshie8 (ganeshie8):

which is same as \[\left\lfloor\dfrac{n}{p^k} \right\rfloor \ge \left\lfloor\dfrac{r}{p^k}\right\rfloor + \left\lfloor\dfrac{n-r}{p^k}\right\rfloor \]

ganeshie8 (ganeshie8):

wid me so far ?

OpenStudy (anonymous):

yess

ganeshie8 (ganeshie8):

recall legendre formula left hand represents the highest exponent of prime \(p\) in \(n!\) , yes ?

OpenStudy (anonymous):

could u type it ? i dont memorize it xD

OpenStudy (anonymous):

wait yes i agree

OpenStudy (conqueror):

Are you and @ganeshie8 going to keep doing this or something ?

OpenStudy (anonymous):

yes yes

ganeshie8 (ganeshie8):

let me ask you a side question whats the highest exponent of "7" in 20! ?

OpenStudy (anonymous):

[20/7]=2

ganeshie8 (ganeshie8):

OK i see you still remember ;)

OpenStudy (anonymous):

yeah like went through my memory xD but 7 seems easy

ganeshie8 (ganeshie8):

we simply interpret the previous inequality for number of primes in the factoriol and conclude our proof

ganeshie8 (ganeshie8):

let me do it in one reply

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@Conqueror ??

ganeshie8 (ganeshie8):

\[\left\lfloor\dfrac{n}{p^k} \right\rfloor \ge \left\lfloor\dfrac{r}{p^k}\right\rfloor + \left\lfloor\dfrac{n-r}{p^k}\right\rfloor \] Left hand side gives the highest exponent of \(p\) in \(n!\) Right hand side gives the highest expone of \(p\) in \(r!(n-r)!\) So the inequality says that the exponent of \(p\) in \(n!\) is greater than the exponent of \(p\) in \(r!(n-r)!\) this is true for all \(p\) consequently \( r!(n-r)! ~| ~n!\)

ganeshie8 (ganeshie8):

\(\blacksquare \)

OpenStudy (anonymous):

wow ! never seen this proof before ! loved it

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

i was thinking of this \(\Large \text{ you can see that }\\ \Large \frac{n!}{(n-r)!}=n(n-1)(n-2)(n-3)\dots (n-r+1)\\ \Large \text{ so we need to show }\\\Large r! |n(n-1)(n-2)(n-3)\dots (n-r+1)\\ \Large \text{W.L.O.G assume } n=r \mod r \text{ thus}\\ \Large n(n-1)(n-2)(n-3)\dots (n-r+1)\equiv r(r-1)(r-2)\dots 2.1 \mod r\\ \Large n(n-1)(n-2)(n-3)\dots (n-r+1)\equiv r! \mod r\\ \Large n(n-1)(n-2)(n-3)\dots (n-r+1)=mr+r!\ \Large \)

OpenStudy (anonymous):

hehe but i wasn't able to end it xD

ganeshie8 (ganeshie8):

we can use this fact : any set of "r" consecutive integers form a complete set of residues "mod r"

OpenStudy (anonymous):

yes this is what i have done

ganeshie8 (ganeshie8):

so one of them will be congruent to 0 mod r making the entire product vanish

ganeshie8 (ganeshie8):

as simple as that !

ganeshie8 (ganeshie8):

what do you mean, you're not able to conclude ?

OpenStudy (anonymous):

vanish w.r to r

OpenStudy (anonymous):

from here :- \(n(n-1)(n-2)(n-3)\dots (n-r+1)\equiv r! \mod r\\ \Large \)

ganeshie8 (ganeshie8):

Consider the product of \(r\) consecutive integers starting from \(x\) : \[x(x+1)(x+2)\cdots (x+r-1)\]

OpenStudy (anonymous):

haha ok ok

ganeshie8 (ganeshie8):

they give you complete set of residues mod r in some order : \[0\cdot 1\cdot 2\cdot 3\cdots (r-1) \pmod r\]

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

that means \(r\) divides the product in question, yes ?

OpenStudy (anonymous):

ok i agree huh

ganeshie8 (ganeshie8):

but you also have (r-1) consecutive integers wid the same reasoning r-1 also divides the product in question, yes ?

OpenStudy (anonymous):

ok got it

ganeshie8 (ganeshie8):

consider 6 consecutive integers : `3,4,5,6,7,8`

ganeshie8 (ganeshie8):

just see that you also have 5 consecutive integers in them so, 5 divides their product

ganeshie8 (ganeshie8):

similarly other integers less than 5 also divide that product

ganeshie8 (ganeshie8):

In general : product of "r" consecutive integers is divisible by r!

ganeshie8 (ganeshie8):

once you know congruences, proofs become short and simple :)

OpenStudy (anonymous):

\( n(n-1)(n-2)(n-3)\dots (n-r+1)\equiv r! \mod r\equiv 0 \mod r\\ n(n-1)(n-2)(n-3)\dots (n-r+1)\equiv (r-1)!\mod r-1 \equiv 0 \mod r-1\\ n(n-1)(n-2)(n-3)\dots (n-r+1)\equiv (r-2)! \mod r-2\equiv 0 \mod r-2\\ n(n-1)(n-2)(n-3)\dots (n-r+1)\equiv (r-3)! \mod r-3\equiv 0 \mod r-3\\ \vdots n(n-1)(n-2)(n-3)\dots (n-r+1)\equiv (3)! \mod 3\equiv 0 \mod 3\\ n(n-1)(n-2)(n-3)\dots (n-r+1)\equiv (2)! \mod 2\equiv 0 \mod 2\\ n(n-1)(n-2)(n-3)\dots (n-r+1)\equiv (1)! \mod 1\equiv 0 \mod 1\\ n(n-1)(n-2)(n-3)\dots (n-r+1)\equiv 0 \mod r!\\ \)

ganeshie8 (ganeshie8):

thats right! but why not keep the proof simple

OpenStudy (anonymous):

hehe :P

ganeshie8 (ganeshie8):

statement : \(r!~ | ~ n(n+1)(n+2)\cdots (n-r-1)\) proof : Clearly the terms in the product represent \(r\) consecutive integers. So they form a complete set of residues in \(\mod r\). That means one of the term in that product is congruent to \(0 \mod r\) making the product divisible by \(r\). Inductively we conclude that the product is divisible by \(r!\) by noticing that the terms in product can form a complete set of resdidues for any integer < \(r\).

ganeshie8 (ganeshie8):

see if that looks good :)

OpenStudy (anonymous):

it is :) btw this is the only proof i have :|

ganeshie8 (ganeshie8):

you also have that combinatoric proof right ?

OpenStudy (anonymous):

combinatoric ?

ganeshie8 (ganeshie8):

\(\large \binom{n}{r}\) is an integer, so \(r! ~|~ n(n-1)(n-2)\cdots (n-r-1)\)

ganeshie8 (ganeshie8):

thats the combinatoric proof ^

ganeshie8 (ganeshie8):

just one line

OpenStudy (anonymous):

haha :P

OpenStudy (anonymous):

ok , i'll force myself to agree :P

ganeshie8 (ganeshie8):

lol almost all combinatoric proofs involve using this binomial coefiicient being integer stuff..

ganeshie8 (ganeshie8):

it is easy to see that the binomial coefficient is integer if you believe in binomial theorem

OpenStudy (anonymous):

I believe that binomial coefficient is triangle integer number then believe in binomial theorem

ganeshie8 (ganeshie8):

not all binomial coefficients are triangular numbers

ganeshie8 (ganeshie8):

triangular numbers are of form n(n+1)/2

OpenStudy (anonymous):

ugh ok

ganeshie8 (ganeshie8):

so you can represent them using binomial coefficient : \[t_n = \binom{n+1}{2}\]

OpenStudy (anonymous):

got it

ganeshie8 (ganeshie8):

ok too much number theory for one day, go prepare for ur exams :)

ganeshie8 (ganeshie8):

i need to understand Hensel's lemma more maybe you can help me after ur exams

OpenStudy (anonymous):

for what u need it ?

OpenStudy (anonymous):

hehe yeah after exams i'll write NT manual solution for burton

ganeshie8 (ganeshie8):

i thought hensel's lemma allows you to solve conruences of type \(r^3+r \equiv a \pmod{p^2}\) ?

OpenStudy (anonymous):

yeah it do

ganeshie8 (ganeshie8):

okay after ur exams teach me then

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i found a funny fact

ganeshie8 (ganeshie8):

?`

OpenStudy (anonymous):

we know that \(\binom{n}{k}=\binom{n-1}{k-1}+\binom{n-1}{k}\)

ganeshie8 (ganeshie8):

pascal's rule

OpenStudy (anonymous):

keep simplifying it until we got it w.r to \(\binom{m_i }{k-k}=1 ,~~~ s.t ~~~ m_i\le n\)

OpenStudy (anonymous):

\(\binom{n}{k} =\) sum of integer combination

ganeshie8 (ganeshie8):

you will reach the first row of paslacl's triangle

OpenStudy (anonymous):

yes yes and simplify more :)

ganeshie8 (ganeshie8):

\[ \sum \limits_{k=0}^n\ \binom{n}{k} = 2^n\] you're not refering to this, right ?

OpenStudy (anonymous):

nope

ganeshie8 (ganeshie8):

can u explain more, i don't get it

OpenStudy (anonymous):

i'll show u

ganeshie8 (ganeshie8):

ok

OpenStudy (anonymous):

\(\begin{align*} \binom{n}{k}&= \binom{n-1}{k}+\binom{n-1}{k-1} \\ &=\binom{n-2}{k-2} +2\binom{n-2}{k-1}+\binom{n-2}{k}\\ &=\binom{n-3}{k-3} +3\binom{n-3}{k-2} +3\binom{n-3}{k-1}+\binom{n-3}{k} \\ &= \binom{n-4}{k-4} +4\binom{n-4}{k-3}+6\binom{n-4}{k-2} +4\binom{n-4}{k-1}+\binom{n-4}{k} \\ &= \dots \text{ so far until we reach all term have } \binom{m_i}{k-k} \text { for some } m_i\le n \end{align*}\)

OpenStudy (anonymous):

so we would have something like this :-

ganeshie8 (ganeshie8):

I know this property, it is called horshow identity or something xD it is easy to see this in pascal triangle

ganeshie8 (ganeshie8):

its called "hockey stick pattern", not horse shoe haha